OpenStudy (anonymous):

ok mulitiply and simplify the rational expressions i do not get this one 18 over x^3 times 5x^2 over 6 i am having problems

OpenStudy (anonymous):

\[18/(x^3) \times (5x^2)/6\] Recall that when you multiply fractions you multiply the numerators(tops) together and the denominators(bottoms) together: \[= \frac{ (18)(5x^2) }{ (x^3)(6) }\] Does that help?

OpenStudy (anonymous):

not really i get that part i am just having problems stying to solve it

OpenStudy (anonymous):

So the next step is to multiply the coefficients: \[= \frac{ 90x^2 }{ 6x^3 }\] (Since 90 is 18 times 5) Do you see how I did that?

OpenStudy (anonymous):

Note that you can't actually solve for x, because this is only an expression(one side of an equation). You are trying to simplify the fraction.

OpenStudy (anonymous):

yes i get it not thsnk you

OpenStudy (anonymous):

But you're not quite done. You can still simplify it further. Do you see how?

OpenStudy (anonymous):

yes 6 goes into the 90 15 times so the answer would be 15 x is that right

OpenStudy (anonymous):

90/6 is 15 yes. but x^2 / x^3 is 1/x So your final answer is 15/x

OpenStudy (anonymous):

when you divide powers with the same base(in this case x) you subtract exponents